# Importance-based Token Merging for Diffusion Models

### v1

@@ 5 Experiments
Our token merging strategy leverages classifier-free guidance for relevance-based token selection, ensuring an efficient and seamless integration without requiring any additional modules or modifications to diffusion models. To showcase its effectiveness, we evaluate our method across three diffusion generation tasks: text-to-image generation, image-conditioned multi-view diffusion, and text-to-video diffusion.
@@ 5.1 Experimental Settings
We use Stable Diffusion 2 (SD) [78] as the base model, a text-to-image latent diffusion model with a U-Net architecture that generates 768x768 images. Our token merging method is compared to ToMeSD [3] and ATC [21]. Due to ATC’s slow inference (several minutes per image), we only display its visual results. For quantitative comparison, we follow previous studies [79, 104, 1, 74] and report FID [24] and CLIP scores [73, 23] for zero-shot image generation on the MS-COCO 2014 validation dataset [49], with 30K randomly sampled image-caption pairs.
@@
We use Zero123++ v1.2 [82] as the base model. Zero123++ is an image-conditioned multi-view latent diffusion model that generates six novel views at a resolution of 320×320. We compare our token merging method to ToMeSD. Following the evaluation protocols of prior work [53, 56], we test on GSO dataset [13], which comprises 30 everyday objects, and compute PSNR, SSIM [98], and LPIPS metrics [108] to evaluate the similarity between the generated images and ground truth. Additionally, we include visual comparisons using in-the-wild images as input to further demonstrate the effectiveness of our method.
@@
We also report TFLOPs, latency, and GPU memory usage, with and without memory-efficient attention (xFormers [40]), on an NVIDIA A5000 GPU. Inference is performed using float16 precision. We estimate TFLOPs for a single diffusion sampling step. For latent diffusion models, the inference cost is measured exclusively in the latent space.
@@
For text-to-image diffusion, we follow the approach used in ToMeSD [3] by merging tokens in the self-attention layers of the first and last blocks of Stable Diffusion. The diffusion process consists of 50 sampling steps, with token pruning applied during the first 6 steps, followed by token merging. The classifier-free guidance (CFG) scale is set to 7.5. For multi-view diffusion, we merge tokens in the self-attention layers of the first two and last two blocks of Zero123++ [82]. The process involves 50 sampling steps, with token pruning applied in the first 10 steps. The classifier-free guidance (CFG) scale is set to 4. For video diffusion, token merging is applied in the first and last blocks of AnimateDiff [17]. The sampling consists of 30 steps, with token pruning during the first 4 steps and a CFG scale of 7.5. To ensure fairness, these settings are consistently applied to both our method and the baselines. The hyper-parameter $p$, which determines the size of the important token pool, is set to 0.4, 0.6, and 0.8 for image, multi-view, and video generation task, respectively. The number of destination tokens ($k$) is consistent across all tasks and follows the setting used in ToMeSD, utilizing 25% of the total tokens.
@@ 5.2 Results
ToMeSD Ours ToMeSD Ours
0 11.88 31.83
0.30 12.20 12.20 31.82 31.82
0.50 13.50 13.42 31.79 31.83
0.60 14.81 14.51 31.80 31.81
0.70 17.46 16.22 31.78 31.79
0.75 20.89 17.75 31.71 31.76
Table 1: Text-to-image generation. We apply ToMeSD [3] and our token merging method to Stable Diffusion [78] across various token merging ratios $r$.
In Tab. 1, we compare different token merging methods applied to Stable Diffusion 2 [78] for text-to-image generation. Our method consistently outperforms ToMeSD [3] in this task, especially at higher token merging ratios ($r$). For example, at $r=0.75$, our method achieves an FID of 17.75 versus ToMeSD’s 20.89. Agglomerative Token Clustering (ATC) [21] is not included due to its prohibitive computational cost, as it is CPU-bound and non-batched, making it impractical for large-scale evaluations. We provide qualitative comparisons in Fig. 5. Note that when the token merging ratio $r$ is small, such as 0.3, our important token pool becomes the whole token set, making it functionally equivalent to ToMeSD.
@@
Table 2: Multi-view diffusion. We apply ToMeSD [3] and our token merging method to Zero123++ [82] with different token merging ratios $r$.
@@
A similar trend can be observed in Tab. 3 and Fig. 7, where we extend the comparison to video diffusion. Across these tests, our method consistently performs better than ToMeSD, both in numerical metrics and in visual quality, particularly in preserving object details. We observed that merging tokens in the temporal layers significantly reduces the generated dynamics. Nonetheless, we present results with token merging applied for both spatial and temporal layers in Tab. 4. In summary, our method outperforms ToMeSD across all three tasks, generating images and videos with better object details.
@@
Inference Costs.
We compare the inference costs of our method and ToMeSD when applied to Stable Diffusion 2 in Tab. 5. As can be seen, both methods show similar improvements in inference times, GPU usage, and TFLOPs. This demonstrates that our token merging strategy can enhance performance without incurring additional computational costs.
@@ 5.3 Ablation Study
0.6 FID $\downarrow$ 14.51 14.56 14.73
CLIP $\uparrow$ 31.81 31.83 31.82
Table 6: Ablation studies on alternative design choices of our method for text-to-image generation. ’w/ top-k dst’ means top-k rather than random selection from the important token pool for destination tokens. ’w/ global ind.’ means independent tokens may also be outside the important token pool, instead of solely within it.
@@
We conduct ablation studies on the text-to-image generation task to examine alternative design choices in our method. As shown in Tab. 6, simply using top-k important tokens as destination tokens leads to worse results. Furthermore, not choosing independent tokens explicitly from the important set (w/ global ind.), leads to a performance drop.
@@
As described in Sec. 4, the size of the important token pool is calculated as $(1+r)\cdot(1+p)$, where $r$ represents the token merging ratio. In Tab. 7, we present results for text-to-image generation using different values of $p$. Setting $p$ too low results in an overly small pool of important tokens, which increases redundancy when randomly selecting destination tokens from it. Conversely, setting $p$ too high includes too many unimportant tokens in the pool, negatively impacting performance.
@@ 6 Conclusions
In this paper, we propose an importance-based token merging method for diffusion models, which maintains generation quality while reducing inference latency. Leveraging classifier-free guidance as an indicator of token importance, our approach strategically allocates computational resources to regions of high relevance to the input condition, thereby enhancing the fidelity of the generated outputs. Our method achieves state-of-the-art performance across diverse tasks, including text-to-image synthesis, multi-view generation, and video generation, highlighting its effectiveness and versatility. Our approach reveals the surprising effect of applying classifier-free guidance at the token level to facilitate compression with minimal impact on quality.
@@
In Appendix A, we present more qualitative comparisons, results on preserving output structure, and experiments with varying diffusion inference steps. In Appendix B, we provide more details about our experimental settings. In Appendix C, we discuss the limitations of our method. In Appendix D, we provide the prompts used to generate qualitative results. We also include a supplementary video for comparisons on text-to-video generation.
@@ Appendix A Additional Results
We provide additional qualitative comparisons between our token merging method and ToMeSD [3]. Additional visual comparisons for text-to-image generation are shown in Fig. 8, and results for multi-view diffusion are presented in Fig. 9. In the supplementary video, we include comparisons on text-to-video generation, using AnimateDiff [17] as the base diffusion model and a merging ratio of 0.7. Furthermore, we provide visual comparisons between ToMeSD [3] and our method across various merging ratios for text-to-image generation in Fig. 10.
Preserving Structure in Early Timesteps.
@@
In Tab. 8 and Fig. 11, we compare the results of ToMeSD [3] under two scenarios: (1) applying token pruning during the early diffusion steps followed by token merging, and (2) using token merging throughout all diffusion steps. The results show that token pruning in the early steps more effectively preserves the generation layout.
@@
Number of Diffusion Inference Steps
$T$ FID $\downarrow$ CLIP $\uparrow$
ToMeSD Ours ToMeSD Ours
20 18.57 17.03 31.77 31.80
30 17.82 16.51 31.78 31.80
50 17.46 16.22 31.78 31.79
Figure 8: Additional comparison of text-to-image generation. The first row shows results from Stable Diffusion (SD) [78], while the subsequent rows show SD combined with ToMeSD [3] and our token merging method. The token merging ratio is 0.7. Best viewed with zoom-in for clarity.
Figure 9: Additional qualitative comparison of multi-view diffusion. We apply ToMeSD [3] and our token merging method to the multi-view diffusion model. We use Zero123++ [82] as the base model and a merging ratio of 0.6. Best viewed with zoom-in.
Figure 10: We provide an additional comparison between ToMeSD [3] and our method when applied to Stable Diffusion [78] across various merging ratios $r$. For reference, the results of Stable Diffusion without token merging are shown on the left. Best viewed with zoom-in for clarity.
Appendix B Additional Experimental Settings
Metrics.
Details on Multi-view Diffusion.
@@ Appendix C Discussion and Limitations
Our method demonstrates broad applicability across diffusion models. However, step-distilled diffusion models [65] do not explicitly compute classifier-free guidance, as this information is already distilled into the final model. An interesting direction for future research could involve refining the distillation process to enable the model to predict an additional output: a classifier-free guidance map, which could then be used for better token merging or other innovative applications.
@@ Appendix D Prompts
Text prompts corresponding to the text-to-image generations in Fig. 5:
@@
  • Elegant teacup with a delicate floral pattern
  • Royal castle with golden towers at sunrise
  • Young musician playing guitar on stage
Text prompts corresponding to the text-to-video generations in Fig. 7:

### v2

@@ 4.1 Experimental Settings
We use Stable Diffusion 2 (SD) [83] as the base model, a text-to-image latent diffusion model with a U-Net architecture that generates 768x768 images. Our token merging method is compared to ToFu [37], ToMeSD [3] and ATC [22]. Due to ATC’s slow inference (several minutes per image), we only display its visual results. For quantitative comparison, we follow previous studies [84, 109, 1, 79] and report FID [25] and CLIP scores [78, 24] for zero-shot image generation on the MS-COCO 2014 validation dataset [52], with 30K randomly sampled image-caption pairs.
We also experiment with a diffusion transformer, PixArt-$\alpha$ [7], for text-to-image generation. We compare our method with ToMeSD and similarly evaluate on the MS-COCO dataset. Unless otherwise stated, “text-to-image” in this paper refers to generation based on Stable Diffusion.
@@
We use Zero123++ v1.2 [87] as the base model. Zero123++ is an image-conditioned multi-view latent diffusion model that generates six novel views at a resolution of 320×320. We compare our method to ToMeSD. Following evaluation protocols of prior work [56, 59], we test on GSO dataset [14], which comprises 30 everyday objects, and compute PSNR, SSIM [103], and LPIPS metrics [113] to evaluate the similarity between the generated images and ground truth. We also include visual comparisons using in-the-wild images as input.
(a) PixArt-$\alpha$ [7](b) ToMeSD [3](c) Ours
Figure 6: Token merging for diffusion transformer. We apply ToMeSD [3] and our method to PixArt-$\alpha$ [7] with a merging ratio of 0.3. Detailed generations are highlighted with red boxes. Best viewed with zoom-in. Please see the supplementary for prompts.
Figure 7: Qualitative comparison of multi-view diffusion. We apply ToMeSD [3] and our token merging method to the multi-view diffusion model. We use Zero123++ [87] as the base model and a merging ratio of 0.6. Our method outputs finer details, as highlighted in red boxes. Best viewed with zoom-in. Please refer to the supplementary for input images.
@@
We report TFLOPs, latency, and GPU memory usage, with and without memory-efficient attention [42], on an NVIDIA A5000 GPU. Inference uses float16 precision. We estimate TFLOPs for a single diffusion sampling step. For latent diffusion models, the inference cost is measured exclusively in the latent space.
@@
For text-to-image synthesis with Stable Diffusion [83], we merge tokens in the self-attention layers of the first and last model blocks, similar to ToMeSD [3]. For PixArt-$\alpha$ [7], we apply token merging to self-attention and cross-attention in the middle half of the model layers (7–20 out of 28) for simplicity. For multi-view diffusion, we merge tokens in the self-attention layers of the first two and last two blocks of Zero123++ [87]. For video diffusion, we merge tokens in the first and last blocks of AnimateDiff [18]. The hyper-parameter $p$, which determines the size of our important token pool, is set to 0.4, 0.6, and 0.8 for image, multi-view, and video generation tasks, respectively. The number of destination tokens ($k$) remains consistent across all tasks and follows the setting used in ToMeSD, utilizing 25% of the total tokens.
@@ 4.2 Results
In Tab. 1 and Fig. 5, we compare different token merging methods applied to Stable Diffusion 2 [83] for text-to-image generation. Our method consistently outperforms baselines, especially at higher token merging ratios ($r$). For example, at $r=0.75$, our method achieves an FID of 17.75 versus ToMeSD [3]’s 20.89. Agglomerative Token Clustering (ATC) [22] is not included due to its prohibitive computational cost, as it is CPU-bound and non-batched, making it impractical for large-scale evaluations. Notably, our method also performs well when used with cross-attention maps (Sec. 4.3). When the merging ratio is small, such as 0.3, our important token pool becomes the whole token set, making it functionally equivalent to ToMeSD.
@@
ToFu ToMe. Ours ToFu ToMe. Ours
0 - - 11.88 - - 31.83
0.30 13.48 12.20 12.20 31.82 31.82 31.82
0.50 15.81 13.50 13.42 31.81 31.79 31.83
0.60 17.32 14.81 14.51 31.81 31.80 31.81
0.70 18.94 17.46 16.22 31.78 31.78 31.79
0.75 19.29 20.89 17.75 31.76 31.71 31.76
Table 1: Text-to-image generation. We apply ToFu [37], ToMeSD [3] and our token merging method to Stable Diffusion [83] across various token merging ratios $r$.
$r$ FID $\downarrow$ CLIP $\uparrow$ Latency (s) $\downarrow$
ToMe. Ours ToMe. Ours
0 27.51 31.30 9.14
0.3 34.23 28.50 30.76 31.05 8.96
0.5 65.46 34.02 29.85 30.68 7.54
0.7 95.46 50.76 28.67 29.93 7.14
Table 2: Comparison of our method with ToMeSD [3] when applied to the diffusion transformer PixArt-$\alpha$ [7].
@@
Table 3: Multi-view diffusion. We compare ToMeSD [3] with our token merging method when applied to Zero123++ [87].
In Tab. 2 and Fig. 6, we show our method significantly outperforms ToMeSD when applied to a diffusion transformer, achieving an FID improvement of 17–48%. This highlights the generalizability of our approach.
@@
A similar trend can be observed in Tab. 4 and Fig. 8, where we extend the comparison to video diffusion. Across these tests, our method consistently performs better than ToMeSD, both in numerical metrics and in visual quality, particularly in preserving object details. We observed that merging tokens in the temporal layers significantly reduces the generated dynamics. Nonetheless, we present results with token merging applied for both spatial and temporal layers in Tab. 5.
@@
Inference Costs.
We compare the inference costs of our method and ToMeSD when applied to Stable Diffusion 2 in Tab. 6. As can be seen, both methods show similar improvements in inference times, GPU usage, and TFLOPs. This demonstrates that our token merging strategy can enhance performance without incurring additional costs.
@@ 4.3 Ablation Study
We conduct ablation studies on the text-to-image generation task to examine alternative design choices in our method. As shown in Tab. 7, simply using top-k important tokens as destination tokens leads to worse results. Furthermore, not choosing independent tokens exclusively from the important set (w/ global ind.), leads to a performance drop.
@@
0.7 FID $\downarrow$ 16.22 16.29 16.43
CLIP $\uparrow$ 31.79 31.79 31.80
Table 7: Ablation studies of our method for text-to-image generation. ’w/ top-k dst’ means top-k rather than random selection from the important token pool for destination tokens. ’w/ global ind.’ means independent tokens may also be outside the important token pool, instead of solely within it.
$r$ FID $\downarrow$ CLIP $\uparrow$
ToMe Ours Ours ToMe Ours Ours
(CA) (CFG) (CA) (CFG)
0.30 12.20 12.17 12.20 31.82 31.82 31.82
0.50 13.50 13.38 13.42 31.79 31.82 31.83
0.70 17.46 16.79 16.22 31.78 31.79 31.79
0.75 20.89 18.17 17.75 31.71 31.75 31.76
Table 8: Comparison of token merging methods applied to Stable Diffusion [83]. CA and CFG denote cross-attention and classifier-free guidance as importance signals, respectively.
@@
Cross-Attention Maps for Token Importance.
In principle, our method can be used with any per-token importance scores. As shown in Tab. 8, we use our method with cross-attention maps instead of CFG. While this approach may require more memory compared to CFG, it remains a strong alternative and highlights the generalizability of our method.
@@
Our important token pool size is $(1-r)\cdot(1+p)$, where $r$ is the token merging ratio. Tab. 9 shows that our method remains robust to the choice of $p$, provided that $p$ is within a reasonable range.
@@ 5 Conclusions
We propose an importance-based token merging method for generation tasks, which maintains generation quality while reducing inference latency. We utilize token importance to strategically allocate computational resources to regions of high relevance to the input condition, thereby enhancing the fidelity of the generated outputs. This novel, simple, and intuitive strategy accelerates various models for free with no modifications needed. Notably, we identify classifier-free guidance as an effective token importance indicator. Our method achieves state-of-the-art performance across diverse tasks, including text-to-image synthesis, multi-view generation, and video generation, highlighting its effectiveness and versatility.
@@
In Appendix A, we present more qualitative comparisons, empirical evidence on the relationship between CFG and token importance, results on consistency models, results on combining our method with orthogonal diffusion acceleration techniques, and experiments with varying diffusion inference steps. In Appendix B, we provide more details about our experimental settings. In Appendix C, we discuss the limitations of our method. In Appendix D, we provide the prompts used to generate qualitative results. We also include a supplementary video for comparisons on text-to-video generation.
@@ Appendix A Additional Results
In Fig. 9, we provide additional qualitative comparisons between ToFu [37], ToMeSD [3], our token merging method, and the variant of our method using cross-attention maps as importance signals. Additional visual comparisons for token merging applied to diffusion transformer are shown in  Fig. 10. Additional results for multi-view diffusion are presented in  Fig. 11. In the supplementary video, we include comparisons on text-to-video generation, using AnimateDiff [18] as the base diffusion model and a merging ratio of 0.7. Furthermore, we provide visual comparisons between ToMeSD [3] and our method across various merging ratios for text-to-image generation in Fig. 12.
Token Importance via Classifier-Free Guidance.
The absolute value of classifier-free guidance (CFG) can be interpreted as a token-level saliency measure, highlighting the tokens that play a crucial role in steering the output toward the given prompt or condition. Empirically, as shown in Tab. 10, removing the top 30% of high-CFG tokens leads to a significant degradation in generation results, whereas removing the bottom 30% has little impact.
No pruning Top 30% Bottom 30%
FID $\downarrow$ 11.88 15.48 12.78
CLIP $\uparrow$ 31.83 31.58 31.88
Table 10: Comparison of pruning (dropping) the top 30% of tokens with the highest CFG values versus the bottom 30% during text-to-image generation using Stable Diffusion [83].
Results on the Consistency Model.
Consistency models [95, 64, 65] typically distill classifier-free guidance (CFG), making the explicit guidance term inaccessible, as they approximate the final guided noise function within a single forward pass. However, our method is not limited to CFG and can leverage any reliable per-token importance signal. As shown in Tab. 11, our token merging, using cross-attention maps as importance signals, remains effective and outperforms the baseline model when applied to the latent consistency model [65].
$r$ FID $\downarrow$ CLIP $\uparrow$ Time Mem.
ToMe. Ours ToMe. Ours (s) $\downarrow$ (GB)$\downarrow$
0 25.38 31.05 0.65 6.75
0.30 25.63 25.63 31.03 31.03 0.60 4.22
0.50 27.80 27.51 30.98 30.99 0.51 3.56
0.60 30.05 29.61 30.90 30.92 0.50 3.37
0.70 35.01 32.49 30.59 30.77 0.48 3.21
0.75 43.28 36.66 30.38 30.60 0.46 3.12
Table 11: Results on the Consistency Model. We show the comparison of token merging methods applied to a 4-step latent consistency model (LCM_Dreamshaper_v7) [65] across various merging ratios $r$.
Combination with Orthogonal Diffusion Acceleration Methods.
In Tab. 12, we demonstrate that our method can be combined with orthogonal diffusion acceleration methods [67, 45, 90] to further accelerate inference while preserving generation quality.
FID $\downarrow$ CLIP $\uparrow$ Time Mem.
(s) $\downarrow$ (GB) $\downarrow$
Ours 16.22 31.79 5.8 3.55
+ DeepCache [67] 15.46 31.81 2.6 3.56
+ FasterDiff. [45] 12.48 31.80 4.2 6.33
+ FRDiff [90] 15.25 31.83 3.5 3.59
Table 12: Combination with Diffusion Acceleration Methods. We show text-to-image generation results by integrating our method with orthogonal diffusion acceleration techniques, using Stable Diffusion [83] as the base model and a merging ratio of 0.7.
Number of Diffusion Inference Steps
$T$ FID $\downarrow$ CLIP $\uparrow$
ToMeSD Ours ToMeSD Ours
20 18.57 17.03 31.77 31.80
30 17.82 16.51 31.78 31.80
50 17.46 16.22 31.78 31.79
Appendix B Additional Experimental Settings
Metrics.
Additional Implementation Details.
For text-to-image generation using Stable Diffusion [83], the diffusion process consists of 50 sampling steps, with the CFG scale set to 7.5. For PixArt-$\alpha$ [7], we perform diffusion sampling for 20 steps, with a CFG scale of 4.5. When computing token similarity for token merging in PixArt-$\alpha$, we find that using pixel location distance of tokens yields better results than feature similarity, and we adopt this approach. For multi-view diffusion using Zero123++ [87], the process involves 50 sampling steps, with the CFG scale set to 4. For video diffusion using AnimateDiff [18], the sampling consists of 30 steps, with a CFG scale of 7.5. To ensure fairness, these settings are consistently applied to both our method and the baselines. For the ablation study investigating the use of cross-attention maps as importance signals, we utilize the averaged attention map from the final model block in Stable Diffusion.
Details on Multi-view Diffusion.
Preserving Structure in Early Time-steps.
Prior work [37] suggests that token pruning (directly dropping tokens) could help preserve structural details. Building on this observation, we found that incorporate token pruning during the early diffusion steps, followed by token merging in later steps, improves generation results. Specifically, we apply token pruning during the first 6, 10, and 4 diffusion inference steps for image, multi-view, and video generation, respectively. Furthermore, the low token variance in the early steps [98] reduces the effectiveness of classifier-free guidance in identifying important tokens. To mitigate this, during these initial steps that involve token pruning, we randomly select one token from each $2\times 2$ region of the feature map as the destination token.
@@ Appendix B Additional Experimental Settings
In Tab. 14 and Fig. 13, we compare the results of ToMeSD [3] under two scenarios: (1) applying token pruning during the early diffusion steps followed by token merging, and (2) using token merging throughout all diffusion steps. The results show that token pruning in the early steps more effectively preserves the generation layout. In Tab. 15, we show that pruning tokens during the first 5-20% denoising steps consistently improves performance across different tasks, highlighting its robustness.
@@
Metric Pruning Steps (m%)
0% 5% 10% 20%
Image
   FID $\downarrow$ 16.27 16.28 16.22 16.13
   CLIP $\uparrow$ 31.77 31.79 31.79 31.79
Multi-View
   PSNR $\uparrow$ 14.73 14.95 14.75 14.80
   SSIM $\uparrow$ 0.783 0.782 0.787 0.785
   LPIPS $\downarrow$ 0.279 0.269 0.268 0.274
Video
   Score $\uparrow$ 79.36 79.59 79.63 79.66
Table 15: We apply token pruning to the first $m\%$ denoising steps (merging ratio = 0.7) and evaluate on three generation tasks.
@@ Appendix C Discussion and Limitations
Our method demonstrates broad applicability across diffusion models. For multi-guidance scenarios (e.g., InstructPix2Pix [5]), weighted averaging of importance signals based on user preferences could be beneficial. For step-distilled diffusion models [70], which already distills classifier-free guidance into the final model, our approach can still be applied by utilizing alternative importance signals, such as attention maps. However, an interesting direction for future research could involve refining the distillation process to enable the model to predict an additional output: a classifier-free guidance map, which could then be used for better token merging or other innovative applications.
@@ Appendix D Prompts
Text prompts corresponding to the text-to-image generations in Figure 5 of the main paper:
  • Elegant teacup with a delicate floral pattern
  • Young musician playing guitar on stage
@@
Text prompts corresponding to the text-to-image generations in Figure 6 of the main paper:
  • A cute cat
  • A real beautiful face
  • A small cactus with a happy face in the Sahara desert
Text prompts corresponding to the text-to-video generations in Figure 8 of the main paper:
@@
In Fig. 14, we show image prompts corresponding to the image-conditioned multi-view generations in Figure 7 of the main paper.
@@
Figure 9: Additional comparison of text-to-image generation. The first row shows results from Stable Diffusion (SD) [83], while the subsequent rows show SD combined with ToFu [37], ToMeSD [3], our method using cross-attention (CA) map, and our method using classifier-free guidance. The token merging ratio is 0.7. Our method outputs finer details, as highlighted in red boxes. Notably, the variant of our method that utilizes the cross-attention map also achieves better generation details compared to baseline methods, demonstrating the generalization ability of our method. Best viewed with zoom-in for clarity.
(a) PixArt-$\alpha$(b) ToMeSD(c) Ours
Figure 10: Additional comparison for token merging applied to diffusion transformer. We apply ToMeSD [3] and our token merging method to PixArt-$\alpha$ [7] for text-to-image synthesis, using a token merging ratio of 0.4. We highlight our generation details with red boxes. Best viewed with zoom-in for clarity.
Figure 11: Additional qualitative comparison of multi-view diffusion. Token merging is applied to the multi-view diffusion model, Zero123++ [87], with merging ratio as 0.6. Our method outputs finer details, as highlighted in red boxes. Best viewed with zoom-in.
Figure 12: We provide an additional comparison between ToMeSD [3] and our method when applied to Stable Diffusion [83] across various merging ratios $r$. For reference, the results of Stable Diffusion without token merging are shown on the left. Our method outputs finer details, as highlighted in red boxes. Best viewed with zoom-in for clarity.
Figure 14: Image prompts for Figure 7 of the main paper.